a quotient of a number and seven is the same as eleven less than four times the numbe. What is the number?
Answer by mananth(16949)

let number be x
x/7 = 4x-11
x=7(4x-11)
x=28x-77
x-28x=-77
-27x=-77
x=-77/-27
x=77/27
....
m.ananth@hotmail.ca